OpenStudy (anonymous):

A rectangle has a perimeter of 20m, express the area of the rect. as a function. of the length of one of its side?

OpenStudy (mayankdevnani):

\[\large \bf Area~of~Rectangle=length \times breadth\] \[\large \bf 20=l \times b\] where,l=length and b=breadth. So,function of l :- \[\large \bf \frac{20}{b}=l\]

OpenStudy (mayankdevnani):

understood? @briah123

OpenStudy (anonymous):

A plot and solution calculated and plotted with Mathematica 9 is attached. This version has the proper line numbering and thank you for the medal.